# Quantum Monte Carlo simulation of light and superlight bipolarons in extended Hubbard–Holstein models on face- and body-centered-cubic lattices

### v1

@@ Abstract
Superlight pairing of bipolarons driven by electron-phonon interactions (EPIs) in face-center-cubic (FCC) and body-center-cubic (BCC) lattices is investigated using a continuous-time path-integral quantum Monte Carlo (QMC) algorithm. The EPIs are of the Holstein and extended Holstein types, and a Hubbard interaction is also included. The number of phonons associated with the bipolaron, inverse mass, and radius are calculated and used to construct a phase diagram for bipolaron pairing (identifying the regions of pairing into intersite bipolarons and onsite bipolarons). From the inverse mass it is determined that for the extended interaction, there is a region of light pairing associated with intersite bipolarons formed in both BCC and FCC lattices. Intersite bipolarons in the extended model at large phonon frequency and large Coulomb repulsion become superlight due to first order hopping effects. The transition temperature of Bose–Einstein condensates of these pairs is estimated. It is determined that intersite bipolarons are associated with regions of high transition temperatures.
@@ 1 Introduction
The discovery of cuprate superconductors stimulated searches for other materials that could superconduct at high temperatures. Among them are the A3C60 materials [1], which are doped from C60 insulators [2] (where A is Cs, Rb or K) and are unconventional superconductors with high transition temperatures (up to 38 K [3]), large Coulomb repulsion, and Jahn–Teller coupling (which has similarities to other electron-phonon couplings). The cesium-doped solid (Cs3C60) can have both BCC (cubic A15) and FCC structures which superconduct at similar temperatures under pressure [3, 4, 5]. The materials have an s-wave order parameter with signatures of unconventional superconductors such as the proximity of Mott and superconducting states [6] and BCS-like theories may be insufficient to explain the superconducting mechanism in these compounds [7, 8, 9]. As such, we are motivated to determine the properties of BCC and FCC bipolarons in systems with strong interactions between phonon modes and electrons in the presence of strong Coulomb repulsion.
We will consider the extended Hubbard–Holstein model (EHHM), which has the following form:
@@
where $g_{ij}$ is a dimensionless electron phonon interaction strength, $c_{i}$ ($c^{\dagger}_{i}$) annihilate (create) electrons and the $a_{j}$ ($a^{\dagger}_{j}$) operators do the same for phonons on site $j$. $t$ is the hopping between neighboring sites. $U$ is the on site Coulomb repulsion (Hubbard $U$). The phonon frequency is $\omega$. The Holstein model corresponds to $g_{ij}=g\delta_{ij}$. The extended Holstein interaction considered here (following Ref. [10]) only shares phonons between sites on nearest-neighour bonds.
In the limit of large phonon frequency, a Lang–Firsov approximation leads to decoupling of phonons and electrons since the average number of phonons in the wavefunction is zero in that limit (where it takes a large quantity of energy to create a phonon), to obtain,
@@
Thus the effective on-site and inter-site coupling strengths are $U^{\prime}=U-2W\lambda$ and $V^{\prime}=-2W\lambda\Phi_{\rm NN}/\Phi_{00}$.
Bipolarons formed in the Holstein-Hubbard model (HHM) and those formed in a truncated EHHM [10] consist of two polarons that can bind into a stable pair when the exchange of phonons overcomes Coulomb repulsion. These categories of bipolarons differ in the spatial extent of their EPI. As shown schematically in Figure 2, the nature of the EPI in the HHM is site-local (electron located at the vibrating atomic site) whereas the EHHM has a long-range EPI (of near-neighbour type). To our knowledge, there are no numerical studies yet of bipolaron properties in either the BCC or the FCC lattice. We note our previous QMC studies of bipolarons on the chain [11], square lattice [12], triangular lattice [13, 12] and simple cubic lattice [14].
@@
where $m^{\ast\ast}$ is the bipolaron (pair) mass, and $n_{b}$ is their density. As $n_{b}$ increases, $T_{\rm BEC}$ also increases until bipolarons start to overlap, at which point the transition temperature saturates before decreasing. Saturation of $T_{\rm BEC}$ is estimated to occur in the vicinity of close packing. Thus, a close packing transition temperature is defined to be [15]
@@
where the bare electron mass is $m_{0}=\hbar^{2}/(2ta^{2})$, and $\Omega_{p}$ is the bipolaron volume in units of the lattice constant $a$. $\Omega_{p}$ needs to be small enough such that inter-bipolaron scattering does not significantly affect Eq. (5).
@@
As can be appreciated from Eqs. (5) and (6), the critical temperature is inversely proportional to the bipolaron mass. Thus in the context of superconductivity, it is important to understand when bipolarons are heavy and when they may be light. The range of electron-phonon interaction affects $m^{\ast\ast}$ in two crucial ways.
@@
First, it affects the polaron mass (i.e. the mass of the two bipolaron constituents), $m^{\ast}\sim m_{0}\exp{(\gamma W\lambda/\hbar\omega)}$ (in the limit of large $\hbar\omega/W$). For the local interaction of the HHM, the lattice deformation is confined to the site occupied by the fermion and is very localised. When the polaron moves, the entire deformation must move to the neighboring site resulting in an exponentially large $m^{\ast}$, since $\gamma=1$ for the HHM. In the EHHM, the lattice on the nearest site is partially pre-deformed. Additional deformation associated with polaron movement is smaller than in the Holstein case, which leads to $\gamma<1$. Thus, EHHM polarons can be exponentially lighter than HHM polarons. We call such polarons “light” [16].
@@
The second effect of the interaction range is related to bipolaron movement. In the HHM, both polarons occupy the same lattice site in the S0 (on-site singlet) configuration. Bipolaron movement is a second-order process that involves an intermediate state with energy excess equal to the bipolaron binding energy $\Delta$. As a result, the bipolaron mass scales as $m^{\ast\ast}\propto(m^{\ast})^{2}\Delta$ in the antiadiabatic limit. Holstein bipolarons are exponentially heavier than Holstein polarons. This result applies to Holstein models in lattices of all dimensionalities. The same scaling $m^{\ast\ast}\propto(m^{\ast})^{2}\Delta$ applies to extended EHHM bipolarons on the BCC lattice, although in this case the constituent polarons are light and therefore the bipolaron is also light. However, EHHM bipolarons on the FCC lattice may display a qualitatively different behavior. If the Hubbard repulsion is strong enough, the two polarons will be forced to occupy two neighbouring sites in S1 configuration. Due to the geometry of the FCC lattice, the S1 bipolaron can move in the first order in polaron hopping in the large $\omega$ limit, as illustrated in Figure 1. The S1 bipolaron mass scales as $m^{\ast\ast}\propto m^{\ast}$. We call such bipolarons “superlight” [16].
@@
Figure 2: One-dimensional schematic of (a) the Holstein-Hubbard model and (b) the extended Holstein-Hubbard model with near-neighbour EPI introduced in Ref. [10] and studied here. The filled circles, empty circles and dashed oval circles represent the electron Wannier orbitals, lattice ions, and nonzero electron-phonon coupling, respectively. The nearest-neighbour electron sites have overlapping orbitals such that an electron can hop via $t$, and $b$ is the intersite spacing.
The goal of this study is to understand whether the above qualitative picture is maintained as $\omega$ is lowered towards the adiabatic limit using a numerically exact QMC simulation method. The paper is organised as follows: In Section 2 an overview of the quantum Monte Carlo scheme is given, in addition to introducing a new triple update needed when simulating FCC lattices in section 2.1 (further details of this scheme can be found in the Appendix). In Section 3 results from the simulations are presented. Finally, in Section 4 a summary can be found.
@@ 2 Method
In this work, we use a continuous-time path-integral quantum Monte Carlo scheme to simulate the HHM and EHHM (see e.g. Refs. [16, 12, 14]). We place two fermions in a $20\times 20\times 20$ box (BCC and FCC lattices, respectively), and then carry out simulations for $\bar{\beta}=t/k_{B}T=20$ (where $k_{B}$ is Boltzmann’s constant and $T$ the temperature) at phonon frequencies of $\hbar\omega/t=1$ and $\hbar\omega/W=1$. A range of $U$ and $\lambda$ are simulated. We adopt twisted boundary conditions on path ends to estimate the mass. Detailed procedures for the simulation and evaluation of estimators for a range of bipolaron properties have been discussed in earlier papers (see Refs. [11, 12, 14], and references therein). Measurements are made every few Monte Carlo steps.
@@
We use the update scheme summarised in Appendix A for BCC lattices (which is essentially the same set of updates as for simple cubic lattices). For FCC lattices, these updates plus an additional update analogous to the triple update are needed to simulate bipolarons on triangular lattices (see Ref. [16]) is also required.
@@ 2.1 Three-kink update for FCC Lattice
Due to the geometry of the FCC lattice, it is possible that a bound bipolaron returns to its initial configuration after three consecutive nearest-neighbour hops of one of the particles. Hence, an update with this property needs to be in the Monte Carlo scheme when simulating FCC lattices to ensure ergodicity. The update procedure is as follows:
  1. 1.
    One of two paths is selected with equal probability $1/2$.
  2. 2.
    A combination of three kinks that is suitable for this update is determined. A kink is selected from the $12$ possible kinks (the number of nearest-neighbour sites) and we label it $\mathbf{l}_{1}$. Then, we choose two other kinks at random ($\mathbf{l}_{2}$ and $\mathbf{l}_{3}$) such that $\mathbf{l}_{1}+\mathbf{l}_{2}+\mathbf{l}_{3}=0$. Satisfying this condition means that the particle returns to its initial site after three hops. None of these three kinks is an antikink to any other.
  3. 3.
    We choose insertion or removal of the (three) kinks with equal probability 1/2.
  4. 4.
    If insertion is chosen during Step 4, we select the imaginary times $\tau_{1},\tau_{2},\tau_{3}$ for the new kinks with equal probability 1/$\bar{\beta}$ from the interval [0, $\bar{\beta}$). The Metropolis condition is given as:
    $P_{\rm add}(\tau_{1},\tau_{2},\tau_{3})={\rm min}\left\{1,\frac{\bar{\beta}}{N_{\mathbf{l}_{1}}+1}\frac{\bar{\beta}}{N_{\mathbf{l}_{2}}+1}\frac{\bar{\beta}}{N_{\mathbf{l}_{3}}+1}\cdot e^{A(D)-A(C)}\right\}$ (7)
    where $C$ and $D$ are the initial and final configurations, respectively. $A(C)$ and $A(D)$ are the bipolaron actions in both initial and final configurations (see Refs. [11, 12, 14]). $N_{\mathbf{l}_{i}}$ ($i=1,2,3$) is the number of kinks of type $\mathbf{l}_{i}$ before the addition update (i.e. in the initial configuration, $C$).
  5. 5.
    On the other hand, if removal of kinks is selected in Step 4, there are two possibilities. The first is to check whether the kink types $\mathbf{l}_{1},\mathbf{l}_{2}$, and $\mathbf{l}_{3}$ chosen from Step 3 already exist in the initial configuration. If any of those selected kinks does not exist in the current path, the acceptance probability is zero and the update is aborted. Otherwise, provided that there is at least one kink of each type, the removal at imaginary times associated with the selected kinks, $\tau_{1},\tau_{2},\tau_{3}$, is carried out with the acceptance probability:
    $P_{\rm remove}(\tau_{1},\tau_{2},\tau_{3})={\rm min}\left\{1,\frac{N_{l_{1}}}{\bar{\beta}}\frac{N_{l_{2}}}{\bar{\beta}}\frac{N_{l_{3}}}{\bar{\beta}}\cdot e^{A(C)-A(D)}\right\}$ (8)
    where $N_{\mathbf{l}_{i}}$ ($i=1,2,3$) is the number of kinks of type $l_{i}$ before removal.
We note that other 3 kink updates are possible, but the selected update is sufficient to ensure ergodicity.

### v2

# Quantum Monte Carlo Simulation of Bipolaron Superconductivity in Extended Hubbard–Holstein models on Face-Centered-Cubic and Body-Centered-Cubic Lattices

@@ Abstract
We investigate superlight pairing of bipolarons driven by electron-phonon interactions (EPIs) in face-center-cubic (FCC) and body-center-cubic (BCC) lattices using a continuous-time path-integral quantum Monte Carlo (QMC) algorithm. The EPIs are of the Holstein and extended Holstein types, and a Hubbard interaction is also included. Effects of adiabaticity are calculated. The number of phonons associated with the bipolaron, inverse mass, and radius are calculated and used to construct a phase diagram for bipolaron pairing (identifying the regions of pairing into intersite bipolarons and onsite bipolarons). From the inverse mass we determine that for the extended interaction, there is a region of light pairing associated with intersite bipolarons formed in both BCC and FCC lattices. Intersite bipolarons in the extended model at intermediate phonon frequency and large Coulomb repulsion become superlight due to first order hopping effects. We estimate the transition temperature, determining that intersite bipolarons are associated with regions of high transition temperatures.
@@ 1 Introduction
The discovery of cuprate superconductors stimulated searches for other materials that could superconduct at high temperatures. Among them are the A3C60 materials [kratschmer1990solid], which are doped from C60 insulators [forro2001electronic] (where A is Cs, Rb or K) and are unconventional superconductors with high transition temperatures (up to 38 K [ganin2008bulk]), large Coulomb repulsion, and Jahn–Teller coupling (which has similarities to other electron-phonon couplings). The cesium-doped solid (Cs3C60) can have both BCC (cubic A15) and FCC structures which superconduct at similar temperatures under pressure [ganin2008bulk, ganin2010polymorphism, ihara2010]. The materials have an s-wave order parameter with signatures of unconventional superconductors such as the proximity of Mott and superconducting states [alloul2013] and BCS-like theories may be insufficient to explain the superconducting mechanism in these compounds [nomura2016, capone2009, gunnarsson1997_sup_in_fullerides]. As such, we are motivated to determine the properties of BCC and FCC bipolarons in systems with strong interactions between phonon modes and electrons in the presence of strong Coulomb repulsion.
One approach would be to study the simplified Hubbard-Holstein model (HHM) in which both the electron-phonon and Coulomb interactions are confined to a single lattice site [Grzybowski2003, macridin2004, Grzybowski2006, Werner2007, DiFilippis2012, Kurdestany2017, Mendl2017]. HHM bipolarons form at strong electron-phonon couplings and generate lattice deformations with delta-function profiles. Coherent movement of such deformations through the lattice is exponentially suppressed. HHM bipolarons are heavy and cannot produce strong superconductivity. In the more complex but more realistic extended Hubbard-Holstein models (EHHM), the electron-phonon interaction extends beyond one site and the lattice deformation has finite dimensions [Alexandrov1999, boncaehhmbipolaron, hague2006effects]. While collective energy gain from the entire deformation still must be large enough to overcome the Coulomb repulsion, the deformation on individual sites is less than in HHM, which is easier to realize. Moreover, extended lattice deformations propagate more coherently than delta-function deformations, leading to exponentially lighter EHHM bipolarons relative to the HHM case.
In this work we will make quantum Monte Carlo simulations of the EHHM with the following form:
@@
where $g_{ij}$ is a dimensionless electron phonon interaction strength, $c_{i}$ ($c^{\dagger}_{i}$) annihilate (create) electrons and the $a_{j}$ ($a^{\dagger}_{j}$) operators do the same for phonons on site $j$. $t$ is the hopping between neighboring sites. $U$ is the on site Coulomb repulsion (Hubbard $U$). The phonon frequency is $\omega$. The Holstein model corresponds to $g_{ij}=g\delta_{ij}$. The extended Holstein interaction considered here (following Ref. [boncaehhmbipolaron]) only shares phonons between sites on nearest-neighbour bonds.
It can be illustrative to consider the limit of large phonon frequency, where a Lang–Firsov approximation leads to decoupling of phonons and electrons (since the average number of phonons in the wavefunction is zero), to obtain,
@@
Thus the effective on-site and inter-site coupling strengths at large phonon frequency are $U^{\prime}=U-2W\lambda$ and $V^{\prime}=-2W\lambda\Phi_{\rm NN}/\Phi_{00}$.
Bipolarons formed in the Holstein-Hubbard model (HHM) and those formed in a truncated EHHM [boncaehhmbipolaron] consist of two polarons that can bind into a stable pair when the exchange of phonons overcomes Coulomb repulsion. These categories of bipolarons differ in the spatial extent of their EPI. Models of this UVU-V type have been studied in the anti-adiabatic limit [micnas1990, Kornilovitch2024]. As shown schematically in Figure 1, the nature of the EPI in the HHM is site-local (electron located at the vibrating atomic site) whereas the EHHM has a long-range EPI (of near-neighbour type). To our knowledge, there are no numerical studies yet of bipolaron properties in either the BCC or the FCC lattice in the adiabatic limit of low phonon frequency. We note our previous QMC studies of bipolarons on the chain [hague2009bipolarons_sing_trip_chain], square lattice [hague2010light], triangular lattice [hague_staggered_ladder, hague2010light] and simple cubic lattice [davenport2012].
@@
where $m^{\ast\ast}$ is the bipolaron (pair) mass, and $n_{b}$ is their density [alexandrov1994]. As $n_{b}$ increases, $T_{\rm BEC}$ also increases until bipolarons start to overlap, at which point the transition temperature peaks before decreasing. Saturation of $T_{\rm BEC}$ is estimated to occur in the vicinity of close packing. Thus, a close packing transition temperature is defined to be [Ivanov1994, Kornilovitch2015, Kornilovitch2024, adebanjo2024]
@@
where the bare electron mass is $m_{0}=\hbar^{2}/(2ta^{2})$, and $\Omega_{p}$ is the bipolaron volume in units of the lattice constant $a$. $\Omega_{p}$ needs to be small enough such that inter-bipolaron scattering does not significantly affect Eq. (5). We calculate $\Omega_{p}=4\pi R^{\prime 3}/3$, where $R^{\prime}=R$ (if $R>b$) and $R=b$ otherwise (since small bipolarons have hard cores).
Figure 1: One-dimensional schematic of (a) the Holstein-Hubbard model and (b) the extended Holstein-Hubbard model with near-neighbour EPI introduced in Ref. [boncaehhmbipolaron] and studied here. The filled circles, empty circles and dashed oval circles represent the electron Wannier orbitals, lattice ions, and nonzero electron-phonon coupling, respectively. The nearest-neighbour electron sites have overlapping orbitals such that an electron can hop via $t$, and $b$ is the intersite spacing.
@@
We now discuss applicability of the ideal-gas BEC formula, Eqs. (5,6), to fullerene physics and to superconductivity in general. Clearly, Eqs. (5,6) neglect the interaction between bipolarons and as such provide only an estimate for the transition temperature. At the same time, in the well-documented case of superfluid helium-4, which is arguably a strongly interacting system, the $T_{c}$ error introduced by neglecting the interaction is only about 40%. We therefore believe that Eqs. (5,6) are adequate to investigate qualitative trends and relative magnitudes of superconducting effects, which is the focus of this paper. Ideally we would solve the full many-body problem, Eq. (1), at finite electron densities and without approximations, but it is still not feasible. Path-Integral QMC employed here is among the most powerful unbiased methods to study many-body Hamiltonians, but its application to the many-fermion sector of Eq. (1) is severely limited by the sign problem. We also note that in strongly interacting electron-phonon systems, even the two-fermion problem, i.e., the bipolaron, is highly complex and nontrivial. Our approach to superconductivity in this and related papers has been to investigate the two-fermion problem without approximations and then to accept the $40\%$ absolute error in $T_{c}$ which still enables a meaningful analysis of relative effects. In practice, we expect this error to be much lower, since estimates of corrections to $T_{c}$ imply a fractional reduction proportional to $n^{2/3}$ for hard-core bosons and proportional to $n$ for inter-boson interactions [micnas1990, alexandrov1986]. The remarkable feature of Eq. (6) is that it contains only two-fermion properties and immediately provides physical insights whenever they are known.
@@
As can be appreciated from Eqs. (5) and (6), the critical temperature is inversely proportional to the bipolaron mass. Thus in the context of superconductivity, it is important to understand when bipolarons are heavy and when they may be light. The range of electron-phonon interaction affects $m^{\ast\ast}$ in two crucial ways. First, it affects the polaron mass (i.e. the mass of the two bipolaron constituents), $m^{\ast}\sim m_{0}\exp{(\gamma W\lambda/\hbar\omega)}$ (in the limit of large $\hbar\omega/W$). For the local interaction of the HHM, the lattice deformation is confined to the site occupied by the fermion and is very localised. When the polaron moves, the entire deformation must move to the neighboring site resulting in an exponentially large $m^{\ast}\sim m_{0}\exp{(W\lambda/\hbar\omega)}$, since $\gamma=1$ for the HHM. In the EHHM, the lattice on the nearest site is partially pre-deformed. Additional deformation associated with polaron movement is smaller than in the Holstein case, which leads to $\gamma<1$. Thus, EHHM polarons can be exponentially lighter than HHM polarons. We call such polarons “light” [hague2007superlight].
@@
The second effect of the interaction range is related to bipolaron movement. In the HHM, both polarons occupy the same lattice site in the S0 (on-site singlet) configuration. Bipolaron movement is a second-order process that involves an intermediate state with energy excess equal to the bipolaron binding energy $\Delta$. As a result, the bipolaron mass scales as $m^{\ast\ast}\propto(m^{\ast})^{2}\Delta$ in the anti-adiabatic limit. Holstein bipolarons are quadratically heavier than Holstein polarons. This result applies to Holstein models in lattices of all dimensionalities. The same scaling $m^{\ast\ast}\propto(m^{\ast})^{2}\Delta$ applies to extended EHHM bipolarons on the BCC lattice, although in this case the constituent polarons are light and therefore the bipolaron is also light. However, EHHM bipolarons on the FCC lattice may display a qualitatively different behavior. If the Hubbard repulsion is strong enough, the two polarons will be forced to occupy two neighbouring sites in S1 configuration. Due to the geometry of the FCC lattice, the S1 bipolaron can move in the first order in polaron hopping in the large $\omega$ limit, as illustrated in Figure 2. The S1 bipolaron mass scales as $m^{\ast\ast}\propto m^{\ast}$. We call such bipolarons “superlight” [hague2007superlight].
@@
The goal of this study is to understand whether the above qualitative picture is maintained as $\omega$ is lowered towards the adiabatic limit using a numerically exact QMC simulation method. The paper is organised as follows: In Section 2 an overview of the quantum Monte Carlo scheme is given, in addition to introducing a new triple update needed when simulating FCC lattices in Section 2.2. In Section 3 results from the simulations are presented. Finally, in Section 4 a summary can be found.
@@ 2 Method
In this work, we use a continuous-time path-integral quantum Monte Carlo scheme to simulate the HHM and EHHM defined in Eq. (1) (see e.g. Refs. [hague2007superlight, hague2010light, davenport2012]). We place two fermions in a $20\times 20\times 20$ lattice constant box, and then carry out simulations for $\bar{\beta}=t/k_{B}T=20$ (where $k_{B}$ is Boltzmann’s constant and $T$ the temperature) at phonon frequencies corresponding to $\hbar\omega/t=1$ and $\hbar\omega/W=1$. The Hamiltonian is simulated for a range of $U$ and $\lambda$. We adopt twisted boundary conditions on path ends to estimate the mass. Detailed procedures for the simulation and evaluation of estimators for a range of bipolaron properties have been discussed in earlier papers (see Refs. [hague2009bipolarons_sing_trip_chain, hague2010light, davenport2012], and references therein). Measurements are made every few Monte Carlo steps.
@@
For BCC lattices, we use the update scheme summarised in Sec. 2.1 (which is essentially the same set of updates as for simple cubic lattices detailed in Ref. [davenport2012]). For FCC lattices, these updates plus an additional update analogous to the triple update used to simulate bipolarons on triangular lattices (Sec. 2.2) are needed (see Ref. [hague2010light]).
2.1 QMC Update Rules and Weighting Scheme
The QMC method used is a continuous-time path integral approach. The electron paths are continuous in time and inter-site electron hops appear as kinks in the paths. We only studied singlet states, meaning that there is no sign problem, since both direct and exchanged configurations are sign positive. Kink insertion or removal is carried out in pairs to ensure that twisted boundary conditions are respected (i.e. the end configurations of the paths are in the same relative positions up to a translation). Following our established algorithm (see Refs. [davenport2012, hague2009bipolarons_sing_trip_chain, hague2010light]), four binary updates were used: (1) A kink of type $\mathbf{l}$ is added to (or removed from) each of both paths. (2) A kink of type $\mathbf{l}$ and its anti-kink $\mathbf{l}$ are added to (removed from) one of the paths. (3) A kink $\mathbf{l}$ and antikink $\mathbf{l}$ are added to (or removed from) different paths (4) A pair of kinks of type $\mathbf{l}$ is added to (or removed from) one of the paths.
On insertion or removal, kinks can be shifted at larger imaginary times (top shift in direction $\mathbf{l}$) or at earlier imaginary times (bottom shift in direction $\mathbf{l}$). Here, $\mathbf{l}$ is a near-neighbor vector.
In all cases, a primary path (denoted $A$) and kink type are selected with probability $1/2$ and $P_{\mathbf{l}}=1/N_{k}$ where $N_{k}$ is the number of nearest neighbours. The other path is denoted $B$. Where kink insertion is chosen, one (two) imaginary time(s) $\tau$ is (are) selected from a uniform probability density $p(\tau)=1/\beta$. For kink removal, a kink of type $\mathbf{l}$ is chosen from path $X$ with probability $1/N_{X\mathbf{l}}$ where $X$ represents path $A$ or $B$.
In the binary updates, correlated shifts indicate that e.g. both kinks relate to a top shift, whereas anti-correlated shifts that each kink comes with a different shift. Correlation and anti-correlation cases are considered to be different subtypes of the update. In the direct configuration, only updates (1) and (2) are available and both correlated and anti-correlated shifts are allowed. In the exchanged configuration, updates (1) and (2) are available as correlated shifts only and (3) and (4) as anti-correlated shifts only. Thus there are four update subtypes attempted in the direct or exchanged configuration. When the path ends are both on the same site, all eight update subtypes are possible. In general, correlated updates maintain the distance between paths and anti-correlated updates change the distance between paths.
Following the update, the electron-phonon action is calculated (see Ref. [hague2007superlight]),
$A[{\bf r}(\tau)]$ $=$ $\frac{z\lambda\bar{\omega}}{2\Phi_{0}(0,0)}\int_{0}^{\bar{\beta}}\int_{0}^{\bar{\beta}}d\tau d\tau^{\prime}e^{-\bar{\omega}\bar{\beta}/2}\sum_{ij}\Phi_{0}[\mathbf{r}_{i}(\tau),\mathbf{r}_{j}(\tau^{\prime})]\left(e^{\bar{\omega}(\bar{\beta}/2-|\tau-\tau^{\prime}|)}+e^{-\bar{\omega}(\bar{\beta}/2-|\tau-\tau^{\prime}|)}\right)$ (7)
$+\frac{z\lambda\bar{\omega}}{\Phi_{0}(0,0)}\int_{0}^{\bar{\beta}}\int_{0}^{\bar{\beta}}d\tau d\tau^{\prime}e^{-\bar{\omega}\tau}e^{-\bar{\omega}(\bar{\beta}-\tau^{\prime})}\sum_{ij}\left(\Phi_{\Delta\mathbf{r}}[\mathbf{r}_{i}(\tau),\mathbf{r}_{j}(\tau^{\prime})]-\Phi_{0}[\mathbf{r}_{i}(\tau),\mathbf{r}_{j}(\tau^{\prime})]\right)$
$-\frac{U}{2}\int_{0}^{\beta}\delta_{\mathbf{r}_{1}(\tau),\mathbf{r}_{2}(\tau)}\,d\tau$
where $\bar{\omega}=\hbar\omega/t$, $\bar{\beta}=t/k_{B}T$, $\Delta\mathbf{r}$ is the twist in the boundary conditions in time and $z$ the number of nearest neighbours.
The detailed balance equations and Metropolis-Rosenbluth conditions for acceptance and rejection of these updates can be found in Ref. [hague2007superlight]. A new update rule for FCC lattices essential for ergodicity is detailed in the next subsection. If a kink removal update is selected and can’t be carried out because there are insufficient kinks of the correct type to remove, then the update is rejected.
@@ 2.2 Three-kink update for FCC Lattice
Analogously to the case of the triangular lattice [hague2010light], electrons on an FCC lattice can return to their original position through three hops. This means that part of the configuration space for the bipolaron containing paths with odd numbers of kinks is not accessible through binary updates. This is generally true for any Hamiltonian where three kinks can be selected such that, $\mathbf{l}_{1}+\mathbf{l}_{2}+\mathbf{l}_{3}=0$. For simple lattices with nearest neighbor hops only, this includes only the triangular and FCC cases. Hence, at lease one three-kink update must be included in the Monte Carlo scheme when simulating FCC lattices to ensure ergodicity. Several three-kink updates are possible in principle, but in practice, an update introducing three kinks to a single path is sufficient to ensure ergodicity. An example of a three-hop loop is shown in Fig. 3.
Figure 3: Schematic showing a three-hop loop on an FCC lattice.
A set of three kinks is selected with the property $\mathbf{l}_{1}+\mathbf{l}_{2}+\mathbf{l}_{3}=0$ with equal probability from all possible sets. The update then attempts to insert or remove these three kinks with equal probability onto one of the paths. As before, where kink insertion is chosen, their imaginary times, $\tau_{1},\tau_{2}$ and $\tau_{3}$ are selected from a uniform probability density $p(\tau)=1/\beta$. For kink removal, a kink of type $\mathbf{l}$ is chosen with probability $1/N_{A\mathbf{l}}$.
The Metropolis condition for insertions is,
$P_{\rm add}={\rm min}\left\{1,\frac{\bar{\beta}^{3}}{(N_{\mathbf{l}_{1}}+1)(N_{\mathbf{l}_{2}}+1)(N_{\mathbf{l}_{3}}+1)}e^{A_{f}-A_{i}}\right\}$ (8)
and for removals,
$P_{\rm remove}={\rm min}\left\{1,\frac{N_{\mathbf{l}_{1}}N_{\mathbf{l}_{2}}N_{\mathbf{l}_{3}}}{\bar{\beta}^{3}}e^{A_{f}-A_{i}}\right\}$ (9)
unless the path contains no kinks of type $\mathbf{l}_{1},\mathbf{l}_{2}$ or $\mathbf{l}_{3}$, in which case the update is rejected. Here $N_{\mathbf{l}_{1}}$ etc. are the number of kinks in the initial configuration, $A_{i}$ is the initial action and $A_{f}$ the final action. Further details on triple updates can be found in Refs. [ganiyuthesis] and [hague2010light].